# Governments Should Mandate Tiered Anonymity on Social-Media Platforms to Counter Deepfakes and LLM-Driven Mass Misinformation

### v1

@@ 9 Cross-Jurisdictional Implementation and Extraterritorial Reach
Framing tiered anonymity as a rights-preserving model is essential, as discussed in the sections above. It does not eliminate anonymity, but conditions amplification on influence. By applying identity obligations only at high reach levels and protecting vulnerable users – like whistleblowers and journalists – the framework can ensure that accountability scales with power, not participation [86].
Resistance will likely be inevitable, especially from the platforms’ side. They may object to the complexity and cost of implementation or worry about user attrition if stricter identity rules push users to fringe platforms. Likewise, some users may attempt to evade tiering by migrating to less-regulated services.
To address this, enforcement must be both staged and strategic. High-leverage jurisdictions like the EU, UK, and US can drive adoption by linking regulatory compliance to market access. App store requirements, advertising standards, and cross-border data flow agreements can reinforce these incentives [84]. Multilateral coordination can ensure that interoperability standards and privacy safeguards are respected, minimizing the risk of regulatory fragmentation [87].
@@ 10 Conclusion
The proposal rests on three pillars. First, empirical evidence from Reddit demonstrates that volunteer moderators already impose proportional gates – karma thresholds, approval queues, and identity checks – as audience size grows [8, 6, 7]. Second, we outline a technically modest implementation that repurposes existing reach telemetry and friction mechanisms such as rate-limited queues and provenance tagging. Third, we trace a viable regulatory pathway: the EU Digital Services Act [9], the UK Online Safety Act [10], and evolving US jurisprudence [88] already link influence to heightened diligence. Tiered anonymity therefore extends, rather than disrupts, the current legal trajectory.
Adopting this model would re-introduce the social friction that recommender systems have eroded, dampening the incentive and impact of large-scale disinformation while sparing ordinary users from onerous disclosure. Future work can focus on refining threshold calibration, explore privacy-preserving credential systems, and evaluate cross-jurisdictional interoperability.

### v2

@@ 9 Enforcement Challenges and Safeguards
Crucially, strong safeguards must be in place to prevent misuse of this system by authoritarian regimes. A tiered identity framework must be framed as rights-preserving: it does not eliminate anonymity, but conditions large-scale amplification on accountability, while protecting vulnerable users like whistleblowers and activists. Authoritarian governments might otherwise attempt to weaponize identity requirements to unmask and persecute dissidents. We note that China, for example, has required real-name registration for “self-media” accounts over 500,000 followers – a move that prompted some users to shut down their accounts for fear of surveillance [34]. To guard against such abuse, our proposal insists on robust due-process limits on identity disclosure (judicial oversight, high thresholds for requests, and transparent challenge mechanisms) [91]. Stored identity data should be encrypted and access logged, with strict minimization and retention policies. These measures, combined with independent oversight, ensure that prospective identity collection enables responsiveness without repression. The goal is a system where accountability scales with influence, yet privacy and free expression remain protected for the vast majority of users.
Resistance will likely be inevitable, especially from the platforms’ side. They may object to the complexity and cost of implementation or worry about user attrition if stricter identity rules push users to fringe platforms. Likewise, some users may attempt to evade the implementation (as per the example from China above) by migrating to less-regulated services. To address this, enforcement must be both staged and strategic. High-leverage jurisdictions like the EU, UK, and US can drive adoption by linking regulatory compliance to market access (e.g. conditioning app store listings or ad business on tiered verification [33]). As discussed a few paragraphs above, multilateral cooperation is equally critical to avoid a fragmented patchwork of rules.
@@ 10 Conclusion
This model extends, rather than upends, current legal trajectories – for example, the EU’s DSA [61] and UK’s Online Safety Act [62] already link greater reach to greater responsibility, and even in the US recent laws and proposals such as the TAKE IT DOWN Act [76] and Section 230 [75] are trending toward conditional accountability. Adopting tiered anonymity would re-introduce the social friction that recommender systems have eroded, dampening the incentive and impact of large-scale disinformation while sparing ordinary users onerous disclosure.
Future work can focus on refining threshold criteria and developing privacy-preserving credential solutions to further bolster the framework. We are confident that calibrating anonymity to influence offers a viable path to a healthier online ecosystem – one that preserves the liberating dog-on-the-Internet [73] anonymity for millions of users, yet ensures that those who command the attention of millions are no longer shielded by total anonymity at the expense of the public good. Each additional safeguard and iteration will bring us closer to an internet that balances free expression with accountability, and privacy with responsibility, in the age of AI-generated media.